OpenStudy (anonymous):

Multiply sqrt(-2) * sqrt(-16) can someone please help me with this one.

myininaya (myininaya):

\[\sqrt{-2}=i \sqrt{2}\] \[\sqrt{-16}=i \sqrt{16}=i \cdot 4= 4i\]

myininaya (myininaya):

Now multiply

myininaya (myininaya):

@eigenschmeigen a and b >=0 in order for that to be true

OpenStudy (anonymous):

good point. i thought something fishy was going on when i wrote that..

myininaya (myininaya):

\[i \sqrt{2} \times 4 i\]

OpenStudy (anonymous):

so is i√2*4i the final answer?

OpenStudy (anonymous):

remember i x i = -1

OpenStudy (anonymous):

so -1*4i

myininaya (myininaya):

don't forget your sqrt(2) part

myininaya (myininaya):

and that extra i shouldn't be there

OpenStudy (anonymous):

so after the i√2*4i what do i put. can you do step by step and explain to me from start to finish

OpenStudy (anonymous):

\[i \sqrt{2} \times 4i = i \times i \times \sqrt{2} \times 4 = -1 \times 4\sqrt{2} = -4\sqrt{2}\]

myininaya (myininaya):

\[i \cdot \sqrt{2} \cdot 4 \cdot i =i \cdot i \cdot 4 \cdot \sqrt{2} \text{ b/c multiplication is communative }\]

myininaya (myininaya):

lol oops i'm too late

OpenStudy (anonymous):

hehe i was speedy

OpenStudy (anonymous):

you guys are great. It has been a long time for me

OpenStudy (anonymous):

so now can you tell me how you got all this i am kind of lost

myininaya (myininaya):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

ok so what is the final answer for the whole problem ? by the way i like your pic, thats cute how do i put something on mine